BRICS Notes Series, 2003

December 17, 2003

This document is also available as PostScript and DVI.

References

[NS-03-3] Abstract, PDF.

Luca Aceto, Zoltn sik, Willem Jan Fokkink, and Anna Inglfsdttir, editors. Slide Reprints from the Workshop on Process Algebra: Open Problems and Future Directions, PA ’03, (Bologna, Italy, 21–25 July, 2003), November 2003. vi+138.

[NS-03-2] Abstract, PostScript, PDF, DVI.

Luca Aceto. Some of My Favourite Results in Classic Process Algebra. September 2003. 21 pp. To appear in the Bulletin of the EATCS, volume 81, October 2003.

[NS-03-1] Abstract, PDF.

Patrick Cousot, Lisbeth Fajstrup, Eric Goubault, Maurice Herlihy, Kurtz Alexander, Martin Rauen, and Vladimiro Sassone, editors. Preliminary Proceedings of the Workshop on Geometry and Topology in Concurrency Theory, GETCO ’03, (Marseille, France, September 6, 2003), August 2003. vi+54.